OpenStudy (anonymous):

\[\sqrt[3]{y ^{2}}/\sqrt[12]{y ^{7}}\]

OpenStudy (anonymous):

\[y^{2/3}/y^{7/12} = y^{2/3-7/12} = y^{1/12} = \sqrt[12]{y}\]